OpenStudy (anonymous):

Verify the identity: (sin4x-sin2x)/(cos4x-cos2x) = TanX

OpenStudy (mathmath333):

is the question complete

OpenStudy (anonymous):

ha oh yea. = tanx sorry about that

myininaya (myininaya):

i would try to be using the double angle identity

myininaya (myininaya):

until no angles were doubled

myininaya (myininaya):

there might be an easier identity but i cannot remember it i would have to search for it

OpenStudy (anonymous):

so far thats what i've been doing. I just struggle when I pick the wrong one.

myininaya (myininaya):

\[\sin(4x)=\sin(2 \cdot 2x)=2 \cdot \sin(2x) \cos(2x)=2 \cdot 2\sin(x)\cos(x) [\cos^2(x)-\sin^2(x)]\]

myininaya (myininaya):

maybe that will help

myininaya (myininaya):

http://www.purplemath.com/modules/idents.htm also the sum identities might work to

OpenStudy (anonymous):

yea, its a good mixture of choosing between the sum/difference, the double angle or the pythgorean identities, i just suck at choosing which to use and at what case.

myininaya (myininaya):

well if you never seen the sum identity talked about on their page then i would go with double angle identity

myininaya (myininaya):

since you have a ton of angles being doubled and you are suppose to show it is tan(x) which requires no doubling

OpenStudy (anonymous):

haha yea, i just realized the sum to product identities are in this chapter, they just aren't used in the example they refer to.

OpenStudy (anonymous):

found a good explanation here.

myininaya (myininaya):

yeah that was the sum identity they used

myininaya (myininaya):

this is another approach here using double angles (it will be a little longer) \[\frac{\sin(4x)-\sin(2x)}{\cos(4x)-\cos(2x)} \\ =\frac{2\sin(2x)\cos(2x)-\sin(2x)}{\cos^2(2x)-\sin^2(2x)-\cos(2x)}\] I used the double angle on top and on bottom Then you could be the bottom in terms of the trig function cos then you can factor top and bottom then something will cancel

myininaya (myininaya):

\[=\frac{2\sin(2x)\cos(2x)-\sin(2x)}{\cos^2(2x)-(1-\cos^2(2x))-\cos(2x)}\\ \frac{2 \sin(2x)\cos(2x)-\sin(2x)}{\cos^2(2x)-1+\cos^2(2x)-\cos(2x)} \] I used the Pythagorean identity there on bottom the bottom in terms of cosine \[\frac{2\sin(2x)\cos(2x)-\sin(2x)}{2\cos^2(2x)-\cos(2x)-1}\] factor top and bottom and you will see something cancel afterwards

myininaya (myininaya):

\[\frac{\sin(2x)(2\cos(2x)-1)}{2\cos^2(2x)-2\cos(2x)+\cos(2x)-1}\\=\frac{\sin(2x)(2\cos(2x)-1)}{2\cos(2x)[\cos(2x)-1]+1[\cos(2x)-1]} \\=\frac{\sin(2x)(2\cos(2x)-1)}{(\cos(2x)-1)(2\cos(2x)+1)}\] well i lied I didn't see that 2 from before so nothing cancels yet

myininaya (myininaya):

but you can write cos(2x)-1 as -2sin^2(x) and write sin(2x)=2sin(x)cos(x) to see something cancel

myininaya (myininaya):

but yeah the sum identities definitely look like a faster approach